OpenStudy (anonymous):

Find,from the definition,the derivative function f(x)=kx^3 where k is a constant

OpenStudy (anonymous):

y'=df(x)/dx=3kx^2

OpenStudy (anonymous):

\[\lim_{h\rightarrow 0}\frac{(k(x+h))^{3}-kx^{3}}{h} = \frac{kx^{3}+3kx^{2}h+3kxh^{2}+kh^{3}-kx^{3}}{h} \] \[=\frac{kh(3x^{2}+3xh+h^{2})}{h} = 3kx^{2}+ 3xh+h^{2}\] when you make h = 0 you just get 3kx^2

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{k(x+h)^3-kx^3}{h}=k \lim_{h \rightarrow 0}\frac{(x^3+3x^2h+3xh^2+h^3)-x^3}{h}\] \[k \lim_{h \rightarrow 0}\frac{3x^2h+3xh^2+h^3}{h}=k \lim_{h \rightarrow 0}\frac{3x^2+3xh+h^2}{1}=k*3x^2\]

OpenStudy (anonymous):

thanx